OpenStudy (anonymous):

in a parent pea plant with the allele pair Gg what is the probability that one gamete will contain the G allele

OpenStudy (anonymous):

 4 total gametes, so 50% chance that it will have G

OpenStudy (anonymous):

I'm going to solve this more algebraically in the hope that seeing the problem approached from more than one angle will possibly help. It's imperative here to understand what the gamete is to understand this question. I think it's easiest to define zygote and gamete together in this instance. Gametes are haploid(containing one of each chromosome, often abbreviated n). Gametes are eggs and sperm. When two gametes fuse together they produce the zygote. The zygote is diploid(has two of each chromosome, often abbreviated 2n) and grows into the pea plant. With that out of the way, the question becomes simple if the following principle is followed: each gamete receives only one allele(so either allele 1 or allele 2... Here G or g). The answer should follow intuitively, but a slightly more rigorous solution is below. Also note in general, if we want to find the chance(whatever%) the following works out: \[\frac{\text{# choices we care about}{\text{# total choices}} \times 100\% \] If this is the case, we could compute this in the following manner: \[\frac{\text{# allele of interest}}{\text{# total alleles}} \times 100\% \] So for example for GG: \[\frac{2}{2} \times 100\% = 100\% \] For gg: \[\frac{0}{2} \times 100\% = 0\% \]

OpenStudy (anonymous):

First equation should have read: \[\frac{\text{# choices we care about}}{\text{# total choices}} \times 100\% \]